OpenStudy (anonymous):

is the integral of (sec x)^2 from zero to .75pi equal negative 1? Can we integrate over this interval?

OpenStudy (anonymous):

No. It is not -1. \[\text{Limit}\left[\text{Sec}[x]^2,x\to \frac{\pi }{2}\right]\to \text{Infinity} \]Refer to the attached plot.

OpenStudy (anonymous):

Thanks